# Multi-Attribute Steering of Language Models via Targeted Intervention

### v1

@@ 3.3 Avoiding Conflicts
By encouraging the steering vectors to be orthogonal, we ensure that each vector operates in a distinct, complementary direction – see Fig. 2(C). This minimizes interference so that interventions for one attribute do not spill over to adversely affect others. Importantly, because of the large activation space of LLM, which is $d$-dimensional, it is possible for all steering vectors to be orthogonal as long as the number of attributes $T\ll d$.
@@ 4.1 Settings
Figure 3: Comparing MAT-Steer and baselines on HelpSteer dataset Wang et al. (2024c). MAT-Steer consistently demonstrates higher win rates compared to baselines using GPT-4 as a judge.
@@
For generation, we use the HelpSteer dataset (Wang et al., 2024d; Dong et al., 2023), which is designed to align LLM outputs with human-preferred characteristics. Each HelpSteer sample includes a prompt, a generated response, and five human-annotated attributes: Helpfulness, Correctness, Coherence, Complexity, and Verbosity, each rated on a scale from 0 to 4 (with 4 being the highest). Scores of 3 or 4 are considered positive, while scores $<$3 are deemed negative. We sample 500 positive and 500 negative instances per attribute. Model outputs are evaluated by GPT-4, which assigns scores for the five attributes following previous work using LLM-as-a-judge (Zheng et al., 2023; Thakur et al., 2024). We report win rates, where a “win” is recorded if MAT-Steer’s output has a higher average score across attributes than the baseline’s.
@@ 4.2 Main Results
For generation tasks, we evaluate our approach using the HelpSteer dataset (Wang et al., 2024d, c), which is designed to align LLM outputs with human-preferred characteristics such as helpfulness, correctness, coherence, complexity, and verbosity. Here, following Zheng et al. (2023); Thakur et al. (2024), each response is scored by GPT-4 on each attribute, and we compute win rates by comparing the overall average attribute scores. As shown in Fig. 3, MAT-Steer consistently achieves higher win rates compared to all baselines. This not only demonstrates that MAT-Steer enhances the desired attributes (e.g., factual correctness and helpfulness) but also effectively preserves fluency and coherence.
@@ 5 Analysis
In this section, we provide a comprehensive analysis of our method, focusing on the trade-offs in intervention and demonstrating the robustness and generalization capabilities of our approach.
@@
To evaluate the impact of our intervention on text generation fluency, we follow prior work (Pham and Nguyen, 2024) and conduct open-ended generation experiments on TruthfulQA using Llama-3.1-8B. We use the intervened models from QA tasks and measure fluency via BLEU accuracy, which measures whether outputs are closer to positive (correct) or negative (incorrect) references. As shown in Fig. 5, MAT-Steer yields higher BLEU accuracy than LoRA fine-tuning (e.g., 45.97 vs. 43.83 SFT) and ITI (45.97 vs. 41.58), indicating more factually correct and coherent outputs.
@@
In the appendix, we provide extensive additional analyses of MAT-Steer along with more numerical results. In particular, in Fig. 7, we demonstrate how the performance of MAT-Steer scales as the number of attributes increases, with increasing win rates over baselines as more attributes are added. We also study the impact of each of MAT-Steer’s components in Table 2, finding that each one contributes substantially and that sparsity makes the largest difference. Furthermore, we compare various token intervention methods in Table 3, finding that MAT-Steer’s improvements are due to selecting the right tokens (not just fewer tokens). In Appendix C, we present further results, including generalization across different model families (see Table 5 and Table 4) – where MAT-Steer also outperforms all baselines – as well as results showing that our method can be effectively combined with ICL and fine-tuning approaches (see Table 6), further enhancing these methods. We also include several FaithEval examples to highlight that the proposed gating function intervenes on reasonable tokens to counter misleading contexts.
@@ Appendix B Further Analysis of MAT-Steer
Ablation Study of MAT-Steer.
Table 2 presents an ablation study on the key components of MAT-Steer evaluated on the TruthfulQA test set. Starting with the base model, which achieves 49.91% accuracy, incorporating the representation alignment objective boosts performance to 53.82%. Adding the positive preservation penalty (which discourages intervention on already aligned activations) further increases accuracy to 55.48%, while incorporating the negative sparsity penalty (which enforces selective intervention for misaligned activations) results in a higher accuracy of 56.73%. Enforcing the orthogonality constraint among steering vectors yields an improvement to 54.37%. Moreover, we observe that normalization plays a crucial role: an ablated version of MAT-Steer without normalization achieves 59.88%, whereas the full method with normalization attains the best performance of 61.94% accuracy. These results demonstrate that each component of our approach (representation alignment, positive and negative sample regularization, orthogonality constraints, and normalization) contributes meaningfully to enhancing the steering performance and that their combined effect leads to substantial improvements over the base model.
Method TruthfulQA
Llama-3.1-8B 49.91
Llama-3.1-8B + Alignment 53.82
Llama-3.1-8B + Alignment + Pos 55.48
Llama-3.1-8B + Alignment + Sparse 56.73
Llama-3.1-8B + Alignment + Orth 54.37
MAT-Steer w/o Normalization 59.88
MAT-Steer 61.94
@@
Appendix C Addtional Numerical Results

### v2

@@ 3.3 Avoiding Conflicts
By encouraging the steering vectors to be orthogonal, we ensure that each vector operates in a distinct, complementary direction – see Fig. 2(C). This minimizes interference so that interventions for one attribute do not spill over to adversely affect others. Importantly, because of the large activation space of LLM, which is $d$-dimensional, it is possible for all steering vectors to be orthogonal as long as the number of attributes $T\ll d=4096$. We further note that our formulation implements orthogonality as a soft penalty via a differentiable regularization, which encourages (but does not enforce) strict orthogonality. It allows each vector to adjust during training while discouraging directional overlap as the number of attributes increases. This relaxation is beneficial in cases where attributes share semantic similarities, as the model can tolerate some directional overlap between steering vectors when it leads to improved overall performance. [^4] We provide an ablation study on different components of MAT-Steer in Section 5.
@@ 4.1 Settings
Figure 3: Comparing MAT-Steer and baselines on HelpSteer dataset Wang et al. (2024c). MAT-Steer consistently demonstrates higher win rates compared to baselines using GPT-4o as a judge.
@@
For generation, we use the HelpSteer dataset (Wang et al., 2024d; Dong et al., 2023), which is designed to align LLM outputs with human-preferred characteristics. Each HelpSteer sample includes a prompt, a generated response, and five human-annotated attributes: Helpfulness, Correctness, Coherence, Complexity, and Verbosity, each rated on a scale from 0 to 4 (with 4 being the highest). Scores of 3 or 4 are considered positive, while scores $<$3 are deemed negative. We sample 500 positive and 500 negative instances per attribute. Model outputs are evaluated by GPT-4o, which assigns scores for the five attributes following previous work using LLM-as-a-judge (Zheng et al., 2023; Thakur et al., 2024). We report win rates, where a “win” is recorded if MAT-Steer’s output has a higher average score across attributes than the baseline’s.
@@ 4.2 Main Results
For generation tasks, we evaluate our approach using the HelpSteer dataset (Wang et al., 2024d, c), which is designed to align LLM outputs with human-preferred characteristics such as helpfulness, correctness, coherence, complexity, and verbosity. Here, following Zheng et al. (2023); Thakur et al. (2024), each response is scored by GPT-4o on each attribute, and we compute win rates by comparing the overall average attribute scores. As shown in Fig. 3, MAT-Steer consistently achieves higher win rates compared to all baselines. This not only demonstrates that MAT-Steer enhances the desired attributes (e.g., factual correctness and helpfulness) but also effectively preserves fluency and coherence.
@@ 5 Analysis
In this section, we provide a comprehensive analysis of our method, focusing on the internal mechanism, the trade-offs in intervention and demonstrating the robustness and generalization capabilities of our approach. Additionally, we provide an ablation study for different components of MAT-Steer.
Internal Mechanism of MAT-Steer.
To evaluate how MAT-Steer generalizes and adapts to both negative samples (where strong intervention is needed) and positive samples (where minimal or no intervention is needed) for a specific task, we conduct an analysis focused on toxicity. After obtaining steering vectors for Table 1 for Llama-3.1-8B, following the setup in ICV (Liu et al., 2024b), we use the ParaDetox dataset (Logacheva et al., 2022) to examine how samples interact with the toxicity steering vector. We randomly sample 100 toxic samples (requiring intervention) and 100 neutral samples (no intervention expected). For each token in every sample, we record the gating weight for the toxicity vector as well as for other attribute vectors (e.g., truthfulness, bias) and the average number of intervened tokens. In addition, we follow Liu et al. (2024b)’s evaluation method by measuring the percentage of toxic samples that flip to neutral (higher is better) after applying MAT-Steer.
Table 2 shows that MAT-Steer correctly selects the toxicity vector with high gating weight, while keeping unrelated attributes largely inactive (0.61 vs. 0.14), demonstrating the importance of sparsity and orthogonality in ensuring targeted steering. This leads to 86% toxicity decrease in ParaDetox. For neutral samples, the gating weight and the number of intervened tokens remain low across all attributes (0.08 and 0.12), showing MAT-Steer’s ability to preserve aligned outputs without unnecessary intervention.
Metric Toxic Non-Toxic
Avg Gating (Toxicity) 0.61 0.08
Avg Gating (Other Attributes) 0.14 0.12
Toxicity Reduction -86% -
Avg # of Intervened Tokens 3.9 0.6
Table 2: Analysis of MAT-Steer on ParaDetox.
@@
To evaluate the impact of our intervention on text generation fluency, we follow a previous work in activation intervention  (Pham and Nguyen, 2024) and conduct open-ended generation experiments on TruthfulQA using Llama-3.1-8B. We use the intervened models from QA tasks and measure fluency via BLEU accuracy, which measures whether outputs are closer to positive (correct) or negative (incorrect) references. As shown in Fig. 5, MAT-Steer yields higher BLEU accuracy than LoRA fine-tuning (e.g., 45.97 vs. 43.83 SFT) and ITI (45.97 vs. 41.58), indicating more factually correct and coherent outputs.
@@
Ablation Study of MAT-Steer.
Table 3 presents an ablation study on the key components of MAT-Steer evaluated on the TruthfulQA test set. Starting with the base model, which achieves 49.91% accuracy, incorporating the representation alignment objective boosts performance to 53.82%. Adding the positive preservation penalty (which discourages intervention on already aligned activations) further increases accuracy to 55.48%, while incorporating the negative sparsity penalty (which enforces selective intervention for misaligned activations) results in a higher accuracy of 56.73%. Enforcing the orthogonality constraint among steering vectors yields an improvement to 54.37%. Moreover, we observe that normalization plays a crucial role: an ablated version of MAT-Steer without normalization achieves 59.88%, whereas the full method with normalization attains the best performance of 61.94% accuracy. Finally, an ablated version of MAT-Steer without any components in Section 3 leads to a significant drop in performance (e.g., 3.86% drop without positive preservation). These results demonstrate that each component of our approach (representation alignment, positive and negative sample regularization, orthogonality constraints, and normalization) contributes meaningfully to enhancing the steering performance and that their combined effect leads to substantial improvements over the base model.
Method TruthfulQA
Llama-3.1-8B 49.91
Llama-3.1-8B + Alignment 53.82
Llama-3.1-8B + Alignment + Pos 55.48
Llama-3.1-8B + Alignment + Sparse 56.73
Llama-3.1-8B + Alignment + Orth 54.37
MAT-Steer w/o Pos 57.37
MAT-Steer w/o Sparse 58.08
MAT-Steer w/o Orth 59.69
MAT-Steer w/o Normalization 59.88
MAT-Steer 61.94
@@
In the appendix, we provide extensive additional analyses of MAT-Steer along with more numerical results. In particular, in Fig. 7, we demonstrate how the performance of MAT-Steer scales as the number of attributes increases, with increasing win rates over baselines as more attributes are added. Furthermore, we compare various token intervention methods in Table 4, finding that MAT-Steer’s improvements are due to selecting the right tokens (not just fewer tokens). In Appendix C, we present further results, including an additional safety benchmark (HH-RLHF) and a reasoning dataset (OBQA) (see Table 5 and Table 6), and generalization across different model families (see Table 8 and Table 7) – where MAT-Steer also outperforms all baselines – as well as results showing that our method can be effectively combined with ICL and fine-tuning approaches (see Table 9), further enhancing these methods. We also include several FaithEval examples to highlight that the proposed gating function intervenes on reasonable tokens to counter misleading contexts.
@@ 7 Conclusion
Limitations
Our method – like the other intervention and trained baselines we compare it to – may struggle in scenarios where the attributes to be aligned are highly complex or unsteerable (Tan et al., 2024). Moreover, like other steering methods, MAT-Steer relies on a small set of data to create steering vectors; we show that MAT-Steer is more data-efficient than several baseline methods, mitigating this issue. In addition, our evaluations have so far been conducted on a select set of tasks and attributes chosen to be representative of standard steering objectives. We leave expanding our approach to an even wider array of tasks as a direction for future research. Our work aims to mitigate LLM risks like bias and toxicity while maintaining performance on other behaviors like helpfulness. As such, MAT-Steer mitigates some of the risks associated with LLMs; however, like other steering methods, MAT-Steer does not eliminate these risks entirely.
@@ Appendix B Further Analysis of MAT-Steer
Appendix C Additional Numerical Results
Additional Results on Other Tasks.
Since our main experiments focus on QA tasks targeting multiple trustworthiness-related attributes, we have extended our evaluation to the HH-RLHF benchmark (Bai et al., 2022) to assess the generalization of MAT-Steer and baselines. In particular, we use the pretrained vectors from QA tasks in Table 1 and test them on the HH-RLHF test set. The HH-RLHF benchmark involves complex, real-world assistant-style queries, covering open-ended topics that require models to be both helpful and harmless. We use GPT-4o to compare generations from MAT-Steer and three baselines: SFT, ICL, and LITO (the strongest baseline in all benchmarks) on the HH-RLHF test set. The GPT-4o win rates are shown in Table 5. These results show that MAT-Steer outperforms all baselines, demonstrating its ability to generalize beyond in-domain QA tasks.
Comparison GPT-4o Win Rates
MAT-Steer vs. ICL 67.07%
MAT-Steer vs. SFT 62.35%
MAT-Steer vs. LITO 59.12%
Table 5: GPT-4o win rates comparing MAT-Steer against other methods on HH-RLHF.
Additionally, we extend MAT-Steer to an explicitly structured reasoning task by incorporating the OpenBookQA (OBQA) (Mihaylov et al., 2018) dataset into our QA tasks (alongside TruthfulQA, Toxigen, and BBQ). OBQA has been used in prior ITI work, such as NL-ITI (Hoscilowicz et al., 2024) and multiple efforts on LLM reasoning (Dubey et al., 2024; Jiang et al., 2024). We compare MAT-Steer on OBQA against three baselines: ICL, SFT (fine-tune the model checkpoint from 3 datasets with additional OBQA training data), and LITO (the strongest ITI baseline in all benchmarks).
As suggested by Fig. 7, most baselines degrade when the number of tasks or attributes increases; therefore, we report both the accuracy on OBQA and the average performance drop on the original QA datasets (TruthfulQA, Toxigen, and BBQ) to assess multi-task robustness. Table 6 demonstrates that while SFT performs slightly better than MAT-Steer on OBQA, it has a significant performance drop of 2.74% on the other tasks. Additionally, ICL has the smallest performance drop (0.14%) but also has the lowest OBQA accuracy (73.40%), indicating limited overall capability. In contrast, MAT-Steer maintains strong performance across all tasks with only 0.32% drop, highlighting its ability to generalize to new reasoning settings while preserving previously aligned behaviors, a key advantage of our multi-attribute steering framework.
Method OBQA (%) Avg. decrease (%)
ICL 73.40 -0.14
SFT 77.92 -2.74
LITO 74.57 -1.12
MAT-Steer 77.46 -0.32
Table 6: Performance on OBQA and average decrease on 3 original QA tasks for different methods.